# Continuous and Discrete-Time Filters: A Unified Operational Perspective

### v1

@@ Abstract
This paper presents a unified tutorial treatment of continuous-time and discrete-time linear time-invariant systems, emphasizing their shared dynamical structure and the physical constraints that differentiate their realizations. Rather than introducing new mathematical tools, the manuscript revisits foundational concepts—transfer functions, poles and zeros, impulse responses, and stability—from an operational perspective rooted in practical signal processing and circuit implementation. First-order systems are used as a minimal yet expressive framework to illustrate how integration, differentiation, filtering, and delay manifest across the Laplace and Z domains. Particular attention is given to causality, bandwidth limitations, sampling effects, and the approximation errors inherent in discrete-time representations. The goal is to bridge the gap between formal mathematical descriptions and the intuition required for reliable system design in mixed analog–digital environments.
@@ III First-Order Continuous-Time Dynamics
$H_{{1^{st}}}(s)=K\frac{s-s_{z}}{s-s_{p}}$ (16)
which approximates an integrator for $s_{p}\ll s_{z}$ and a derivator when $s_{z}\ll s_{p}$, but only at for frequencies $s\in[s_{z},s_{p}]$ and sufficiently far from the endpoints.
@@ IV From Continuous-Time Signals to Discrete-Time Sequences
Here a link between continuous-time signals $x(t)$ and their discrete-time counterparts $x[n]$ established through sampling. This connection naturally gives rise to the discrete-time Fourier transform (DTFT) and, under periodic sampling, the discrete Fourier transform (DFT). DT systems implement signal-conditioning functions by means of basic arithmetic operations (addition, subtraction, and scaling) on samples of a time-sequence. Since such processing generally requires access to past samples, a delay function is essential. This allows implementing memory as feeding a sample to a chain of delays makes it available to future computations. In digital implementations, this delay is typically provided by clocked registers, each storing a sample for one clock period. Analog sampled‑data systems, on the other hand, often use capacitors to dynamically store analog voltages, with multi‑phase switching to separate sampling and processing phases. Regardless of the implementation domain, a feed‑forward system produces an output that depends only on past inputs:
@@
Conversely, introducing an additional path feeding output samples back to the input makes the current output also depends on its own previous samples, leading to the recursive relationship:
@@
In DT-systems reproducing an impulse is straightforward as it only requires a single non-zero sample in the time-sequence. Conversely in the time-continuous counterpart, it is not feasible to implement an ideal Dirac delta as it requires infinite bandwidth and amplitude. Feed-forward systems, described generally by Eq. 27, produce at most $m$ non-zero output samples in response to an impulse and are therefore classified as finite-impulse-response (FIR) filters. The recursive form given in Eq. 28 yields an impulse response that, in principle, extends indefinitely; systems of this type are therefore classified as infinite‑impulse‑response (IIR) filters. Stability requirements, however, constrain the response to decay exponentially over time. While a decaying exponential only asymptotically approaches zero, practical implementations exhibit effectively finite impulse responses as the signal eventually falls below the resolution limit of the system (whether determined by the finite precision of a digital representation or by the noise floor in an analog one).
@@ IV-A The Z-transform
$X_{\textit{in}}(z)=\mathcal{Z}\{x_{\textit{in}}[n]\}=\sum_{n=0}^{\infty}x_{\textit{in}}[n]\cdot z^{-n}$ (29)
If the sequence is of finite length, there exists an index $m$ such that $x_{\textit{in}}(n)=0$ for all $n\geq m$. In this case, the sum in Eq. 35 contains only a finite number of non-zero terms. The implication is that the function $X_{in}(z)$ is well-defined across the entire complex plane. Essentially, the Z-transform encodes the information about the input signal into a mapping of the complex plane $X_{\textit{in}}(z):\mathcal{C}\rightarrow\mathcal{C}$; alternatively the map can be decomposed into two real-valued functions which provide, for instance, the amplitude and phase components of the complex image. Expressing $z$ in polar form makes explicit the role of the magnitude $\rho$ in geometrically scaling successive samples and the role of the phase $\theta$ in preserving their temporal ordering:
@@
The relationship between complex phase and time‑delay follows directly from the temporal structure of periodic signals. For a sinusoid of angular frequency $\omega$, the instantaneous phase evolves linearly in time: $\theta(t)=\omega t+\theta_{0}$. Using Euler’s formula, the real cosine signal can be expressed as the real part of a complex exponential:
$\mathit{cos}(\omega t+\theta_{0})=\operatorname{Re}\left[e^{j(\omega t+\theta_{0})}\right]$ (31)
This equation highlights how a time delay corresponds to a multiplication by a pure phase in the complex exponential representation:
$\mathit{cos}\left(\omega(t-\Delta t)+\theta_{0}\right)=\operatorname{Re}\left[e^{i(\omega t+\theta_{0})}\cdot e^{-i\omega\Delta t}\right]$ (32)
@@
where the delay between consecutive samples is equal to one sampling period. In the Z-transform of the sequence a delay of $k$ samples corresponds to the phase shift:
$e^{i\Delta\theta}=e^{-i\omega\Delta t}=e^{-i\frac{\omega}{F_{s}}k}$ (34)
where $\omega$ denotes the angular frequency of the signal itself, if periodic, or of its harmonic components in the non-periodic case. This shows how the phase of the factor $z^{-k}$ accounts for the cumulative phase lag due to a delay of $k$ samples.
In order for the Z-transform to be well-defined, the infinite series of Eq.35 needs to converge to a non-trivial value. The convergence critically depends on the properties of the signal. For periodic or infinite-duration signals, the region of convergence (ROC) always excludes the interior of the unit circle ($|z|<1$) because the geometric terms $\rho^{-n}$ grow unboundedly with $n$. This exactly mirrors the stability condition for $s$-domain poles requiring negative real parts. Indeed, the relationship between exponentials and geometric series implies that $\operatorname{Re}(s_{p})\leq 0$ corresponds to $|z_{p}|\leq 1$. Conversely, for finite-length sequences, the infinite series reduces to a polynomial in $z^{-1}$ which converges for all finite values with the exception of the origin. The points at infinity are also excluded from the ROCs because they represent a zero that trivializes the transform.
@@ IV-B Discrete-Time Fourier Transform
Restricting the complex variable to the unit circle corresponds to discarding the amplitude information while retaining only the phase term of the transform, yielding:
$X_{\textit{in}}(z)|_{z=e^{i\theta}}=X_{\textit{in}}(\theta)=\sum_{n=0}^{\infty}x_{\textit{in}}[n]\cdot e^{-i\theta n}$ (35)
Having previously established the relationship between time-delay and phase difference, it is convenient to interpret the index $n$ as a discrete time measured in samples rather than a dimensionless integer. Consequently, the total phase ($\theta\cdot n$) can be factorized in the product of the discrete time and a discrete angular frequency $\omega_{d}$. Trivially the latter corresponds to the complex phase $\theta$ measured in radians per sample. This leads to an alternative formulation of the Z-transform known as the DT Fourier transform (DTFT) of a causal sequence:
$X_{\textit{in}}(\omega_{\mathit{d}})=\sum_{n=0}^{\infty}x_{\textit{in}}[n]\cdot z[n]\large|_{\rho=1}=\sum_{n=0}^{\infty}x_{\textit{in}}[n]\cdot e^{-i\omega_{d}n}$ (36)
where the complex variable reduces to a repeating sequence of samples over the unit circle. Geometrically, the points on the unit circle represent the phase accumulation corresponding to a specific time delay. Keeping track of how the phase varies with time is equivalent to a frequency analysis. In fact, from Eq. 36 choosing a specific frequency of interest, $\omega$, corresponds to fixing the uniform spacing of the samples over the unit circle. This spacing is given by the discrete angular frequency:
$\omega_{d}=\frac{\theta}{1}=\frac{\omega}{F_{s}}$ (37)
@@
The ratio $2\pi/\omega_{d}$ gives the number of samples taken per period of the original signal, $x_{in}(t)$, or of its harmonic component at $\omega$. To uniquely determine a sinusoid, up to an integer multiple the sampling frequency, at least three independent points are needed. This allows the amplitude, frequency and initial phase to be resolved. Having only two equally spaced samples per period produces a repeating sequence that admits an infinite family pf amplitude-phase pairs. With slightly more than two samples per period ($\omega<\pi F_{s}$) the sinusoid is sampled at distinct phase points in each period, enabling unique identification. The amplitude and phase of a discrete spectral component of the signal at a chosen frequency are obtained by evaluating the DTFT according to Eq. 37.
@@
The DTFT in Eq. 36 is a continuous and periodic function of the discrete angular frequency. The complex exponential introduces a periodicity of $2\pi$, causing the frequency axis to repeat. Moreover, the symmetries introduced by the trigonometric functions:
$\begin{cases}e^{i\theta}=e^{i(\theta+\pi\cdot n_{even})}\\[8.0pt] e^{i\theta}=-e^{i(\theta+\pi\cdot n_{odd})}\end{cases}$ (38)
@@
are responsible for the periodicity of the magnitude spectrum which is repeated in the intervals $\pi\cdot n_{even},\pi\cdot(n_{even}+1)]$ and mirrored in the intervals $\pi\cdot n_{odd},\pi\cdot(n_{odd}+1)]$. This aliasing effect restricts the meaningful analysis of discrete‑time signals to the Nyquist range, $\omega_{d}\in[0,\pi]$, corresponding to continuous frequencies $f\in[0,F_{s}/2]$. Beyond this range is not possible to have a unique mapping between time-continuous signals and their sample sequences. The aliasing refers all the compatible signals at higher frequencies that are compatible with a given sequence. For instance, a discrete sinusoid sampled at $F_{s}=\omega/2\pi$ is identical to a constant sequence, as it is sampled once per period always at the same phase point.
@@
$\sum_{n=0}^{\infty}cos(\omega_{0}n)\cdot e^{-i\omega_{d}n}=\pi\sum_{n=-\infty}^{\infty}\left[\delta(\omega_{d}\pm\omega_{0}-2\pi n)\right]$ (39)
@@ IV-C Discrete Fourier Transform
Because the DTFT requires an infinite sequence, practical implementations truncate the signal to a finite length $N$ and compute the discrete Fourier transform (DFT) instead, which is also a sequence of the same length:
$X_{\textit{in}}(\omega_{k})=X_{\textit{in}}[k]=\sum_{n=0}^{N-1}x_{\textit{in}}[n]\cdot e^{-in\omega_{k}}$ (40)
where
@@
which highlights the fact that $\omega_{k}$ is a sampling of the continuous variable $\omega_{d}$. The DFT is therefore a uniform sampling of the DTFT at N points over the interval $0,2\pi F_{s})$. The Nyquist frequency is found at $k=N/2$ and the spectral resolution is given by the distance between consecutive evaluation frequencies which provides the bin width $\Delta\omega=2\pi F_{s}/N$.
Eq. 40 rests on the fundamental fact that any sequence of $N$ equally spaced samples which is also perfectly periodic in time can be exactly reconstructed as a finite linear combination of sinusoids whose frequencies are given by Eq. 41. Equivalently, because discrete complex exponentials form an orthonormal basis for all periodic sequences, the set of frequencies that uniquely represent the signal is entirely determined by the sequence length and the sampling rate. The periodicity of the time sequence is actually enforced by the mathematical duality of sampling and periodic repetition between the time and frequency domain. Namely, performing a sampling of the DTFT implicitly converts the finite sequence into a periodic one with $x_{\textit{in}}[N\cdot n]=x_{\textit{in}}[N\cdot(n+1)]$. Finally since both terms under the sum in Eq. 40 are periodic modulo $N$, the sampled spectrum can also be repeated across the discrete frequency axis.
@@ V Discrete-Time LTI Systems and the Z-Domain
While the transfer function of a continuous‑time system, Eq. 2, is derived via the Laplace transform, discrete‑time systems are naturally characterized using the Z‑transform, leading to a transfer function of the form polynomial ratio of the type:
@@
In the DT domain, the infinitesimal operations of differentiation and integration are replaced by finite differences and accumulation respectively, providing a simple algebraic relation between the samples in the sequence. Nevertheless, Z‑domain transfer functions retain significant advantages: they provide clear insight into frequency‑domain behavior and impulse‑response characteristics, and they simplify the analysis of interactions between subsystems. These features are essential for enabling systematic, modular design of discrete‑time systems. As for the CT case, strict causality imposes a backward definition of the derivatives. Therefore the time-discrete derivatives of a sequence $x[k]=x(t)|_{t=k\cdot T_{s}}$ are defined as:

### v2

@@ Abstract
Continuous-time (CT) and discrete-time (DT) linear time-invariant (LTI) systems are commonly introduced through distinct mathematical formalisms, which can obscure their underlying dynamical equivalence. This tutorial presents a unified treatment of first-order CT and DT systems, emphasizing their shared modal structure and stability properties. Beginning with transfer functions and pole–zero representations in the Laplace domain, canonical first-order low-pass and high-pass dynamics are examined from an operational perspective. The discussion then transitions to discrete-time sequences and the Z-transform, highlighting geometric sequences as eigenfunctions of DT systems and establishing the correspondence between the left half of the s-plane and the interior of the unit circle in the z-plane. Practical discretization and sampled-data implementations are analyzed to illustrate how continuous-time dynamics are reinterpreted through recursion and accumulation in discrete-time realizations. By maintaining structural symmetry between domains, the manuscript consolidates established concepts into a coherent framework linking mathematical representation, physical realizability, and implementation.
@@ III First-Order Continuous-Time Dynamics
$H_{phys}(s)=K\cdot\frac{s-s_{z}}{s-s_{p}}\cdot\frac{1}{s-s_{nd}}$ (16)
Where $s_{nd}$ is the non-dominant pole responsible for properness and which approximates an integrator for $s_{p}\ll s_{z}$ and a derivator when $s_{z}\ll s_{p}$, but only at for frequencies $s\in[s_{z},s_{p}]$ and sufficiently far from the endpoints as depicted in Fig. 1.
Figure 1: Approximated Lapace-domain transfer functions illutrating filter characteristics influenced by parasitic physical elements. Here $s_{p}$ denotes the first-order pole, $s_{z}$ the first order zero, and $s_{nd}$ the additional non-dominant pole that elevates the system to second-order behavior.
@@ IV From Continuous-Time Signals to Discrete-Time Sequences
This section establishes how sampling links CT signals $x(t)$ to DT sequences $x[n]$ and shows how this relationship leads to the discrete-time Fourier transform (DTFT) and, under periodic sampling, to the discrete Fourier transform (DFT).
DT LTI systems implement signal-conditioning functions through the basic arithmetic operations of addition, subtraction, and scaling—applied to samples of a time sequence. To perform DT integration and differentiation, as well as their combination into specific filtering responses, access to past samples is required. This is achieved using delay elements which effectively enables memory: building arbitrary chains of delays grant future computations access to past samples. In digital implementations, this delay is typically realized with clocked registers, each storing a sample for one clock period. Analog sampled-data systems, by contrast, often employ capacitors to store analog voltages dynamically, using multi-phase switching to separate sampling and processing phases. Regardless of the implementation choice, the same mathematical models are applicable. The simpler one is the generic feed‑forward structure where the current output only depends on past inputs:
@@
A richer class of systems is obtained by incorporating feedback, a signal path from the output back to the input, which introduces dependence on past output samples and results in the general recursive sequence:
@@
Unlike the CT case, where an ideal impulse (Dirac delta) is by definition unrealizable,requiring infinite bandwidth and amplitude, its DT counterpart is straightforward to implement, consisting merely of a single non-zero sample in a sequence. Feed-forward systems, described generally by Eq. 27, produce at most $m$ non-zero output samples in response to an impulse and are therefore classified as finite-impulse-response (FIR) filters. The recursive form given in Eq. 28 yields an impulse response that, in principle, extends indefinitely because of feedback; systems of this type are therefore classified as infinite‑impulse‑response (IIR) filters. Stability requirements, however, constrain the response to decay exponentially over time. While a decaying exponential approaches zero only asymptotically, practical implementations exhibit finite responses as the signal eventually falls below the resolution limit of the system (whether determined by the finite precision of a digital representation or by the noise floor in an analog one).
@@ IV-A The Z-transform
$X_{\textit{in}}(z)=\mathcal{Z}\{x_{\textit{in}}[n]\}=\sum_{k=0}^{\infty}x_{\textit{in}}[k]\cdot z^{-k}$ (29)
If the sequence is of finite length, there exists an index $m$ such that $x_{\textit{in}}(n)=0$ for all $n\geq m$. In this case, the sum in Eq. 29 contains only a finite number of non-zero terms. The implication is that the function $X_{in}(z)$ is well-defined across the entire complex plane. Essentially, the Z-transform encodes the information about the input signal into a mapping of the complex plane $X_{\textit{in}}(z):\mathcal{C}\rightarrow\mathcal{C}$. Alternatively, the map can be decomposed into polar components, two real-valued functions which provide the amplitude and phase of the complex image. Expressing $z$ in polar form makes explicit the role of the magnitude $\rho$, which geometrically scales successive samples, and the role of the phase $\theta$, which tracks the temporal ordering:
@@
The relationship between complex phase and time‑delay follows directly from the structure of periodic signals. For a sinusoid of angular frequency $\omega$, the instantaneous phase evolves linearly in time: $\theta(t)=\omega_{0}t+\theta_{0}$. Using Euler’s formula, the cosine function can be written as the real part of a complex exponential:
$\mathit{cos}\left(\omega_{0}(t\pm\Delta t)+\theta_{0}\right)=\operatorname{Re}\left[e^{i(\omega_{0}t+\theta_{0})}\cdot e^{\pm i\omega_{0}\Delta t}\right]$ (31)
This relation, highlights how a time delay becomes equivalent to a phase shift which can be obtained by multiplication of a pure phase in the complex exponential representation.
@@
Here the delay between consecutive samples is by construction equal to one sampling period $T_{s}=1/F_{s}$. In the Z-transform of the sequence the delay information is retained by the factor $z^{-k}$ which accounts for the cumulative phase lag corresponding to a delay of $k$ samples. For a sinusoidal input with angular frequency $\omega_{0}$:
$\angle(z^{-k})=e^{i\Delta\theta_{k}}=e^{-i\omega_{0}\Delta t_{k}}=e^{-i\frac{\omega_{0}}{F_{s}}k}$ (33)
For a generic input, $\omega_{0}$ denotes the angular frequency of the signal itself, if periodic, or of its harmonic components in the non-periodic case.
By contrast, the modulus of the complex variable $\rho^{-k}$ governs the convergence of the series in Eq. (29). For the Z‑transform to be well‑defined, the infinite series must converge to a finite, non‑trivial value. Convergence depends critically on the nature of the signal. For periodic or infinite‑duration signals, the region of convergence (ROC) always excludes the interior of the unit circle ($|z|<1$), because their geometric factors would grow boundless with $k$.
This condition mirrors the stability requirement for ss‑domain poles, which mandates negative real parts. Indeed, the correspondence between exponentials and geometric series implies that $\operatorname{Re}(s_{p})\leq 0$ translates to $|z_{p}|\leq 1$ in the discrete domain.
For finite‑length sequences, the infinite series reduces to a polynomial in $z^{-1}$, which converges for all finite $z$ except the origin. Points at infinity are likewise excluded from the ROC, as they would correspond to a zero that trivializes the transform.
@@ IV-B Discrete-Time Fourier Transform
A frequency-domain representation that provides the spectral information of a signal is obtained from the Z‑transform by restricting the complex variable to the unit circle:
$X_{\text{in}}(z)\big|{z=e^{i\theta}}=X_{\text{in}}(\theta)=\sum_{k=0}^{\infty}x_{\text{in}}[k]\cdot e^{-i\theta k}$ (34)
where the complex variable now traverses a periodic sequence of samples along the unit circle. Geometrically, each point on the unit circle encodes the phase accumulation associated with a given time delay. Tracking how this phase evolves with time is equivalent to performing a frequency analysis. In fact, interpreting the index $k$ as discrete time measured in samples, rather than as a dimensionless integer, establishes the relationship between the complex phase $\theta$ and the discrete angular frequency $\omega_{d}$, expressed in radians per sample:
$\theta=\omega_{d}\cdot 1=\frac{\omega}{F_{s}}$ (35)
where $F_{s}$ is the sampling frequency. Equation (35) formalizes the correspondence between a continuous‑time frequency of interest $\omega$ and the resulting uniform angular spacing of samples on the unit circle.
The ratio $2\pi/\omega_{d}$ then gives the number of samples taken per period of the original continuous‑time signal $x_{\text{in}}(t)$—or, equivalently, per period of its harmonic component at frequency $\omega$.
@@
To uniquely determine a sinusoid, up to an integer multiple the sampling frequency, at least three independent points are needed. This allows the amplitude, frequency and initial phase to be resolved. Having only two equally spaced samples per period produces a repeating sequence that admits an infinite family of amplitude-phase pairs. With slightly more than two samples per period ($\omega<\pi F_{s}$) the sinusoid is sampled at distinct phase points in each period, enabling unique identification.
A discrete sinusoid identifies a discrete spectral component of the transformed signal, whose amplitude and phase values are obtained by evaluating the DTFT of Eq. 34 at the discrete frequency given by Eq. 35.
@@
The DTFT in Eq. 34 is a continuous and periodic function of the discrete angular frequency. The complex exponential introduces a periodicity of $2\pi$ radians, causing the frequency axis to repeat indefinitely. In addition, the symmetries properties of trigonometric functions introduce additional structure. In fact,
$e^{i(\theta+\pi\cdot n)}=(-1)^{n}\cdot e^{i\theta}$ (36)
@@
implies that shifting the frequency by integer multiples of $\pi$ either preserves or flips the sign of the complex exponential depending on the parity of $n$. As a result, the magnitude spectrum repeats identically over intervals $\pi\cdot n_{\text{even}},\pi\cdot(n_{\text{even}}+1)]$ and appears mirrored over intervals $\pi\cdot n_{\text{odd}},\pi\cdot(n_{\text{odd}}+1)]$. This effect is called aliasing and restricts the meaningful analysis of discrete‑time signals to the Nyquist range, $\omega_{d}\in[0,\pi]$, corresponding to continuous frequencies $f\in[0,F_{s}/2]$. Beyond this range, it is no longer possible to establish a unique mapping between a continuous‑time signal and its sampled sequence: multiple higher‑frequency signals become indistinguishable from lower‑frequency ones when sampled at a given rate. For instance, a discrete sinusoid sampled at $F_{s}=\omega/2\pi$ is identical to a constant sequence, as it is sampled once per period always at the same phase point.
@@
$\sum_{k=0}^{\infty}cos(\omega_{0}k)\cdot e^{-i\omega_{d}k}=\pi\sum_{k=-\infty}^{\infty}\left[\delta(\omega_{d}\pm\omega_{0}-2\pi k)\right]$ (37)
@@ IV-C Discrete Fourier Transform
Because the DTFT is defined for an infinite-length sequence, practical implementations must truncate the signal to a finite length $N$ and instead compute the discrete Fourier transform (DFT), which yields a discrete sequence of the same length:
$X_{\textit{in}}(\omega_{k})=\tilde{X}_{\textit{in}}(e^{-in\omega_{k}})=\sum_{k=0}^{N-1}x_{\textit{in}}[k]\cdot e^{-in\omega_{k}}$ (38)
where $\tilde{X}_{in}(z)$ is the finite approximation of the Z-transform and the discrete frequency points are given by:
@@
This formulation highlights that $\omega_{k}$ corresponds to a uniform sampling of the continuous frequency variable  [^1] Although called the discrete frequency, $\omega_{d}$ is in fact a continuous variable: it is simply the continuous angular frequency $\omega$ normalized by the sampling rate $F_{s}$. The descriptor ”discrete” properly applies to the time-domain sequence, not to the frequency variable itself. $\omega_{d}$. The DFT thus represents a uniformly sampled version of the DTFT at $N$ points over the interval $0,2\pi F_{s})$.
The Nyquist frequency corresponds to the index $k=N/2$, and the spectral resolution—determined by the spacing between consecutive frequency samples—is given by the bin width $\Delta\omega=2\pi F_{s}/N$.
The validity of Eq. (38) follows from the fact that any uniformly sampled, perfectly periodic sequence of length $N$ admits an exact decomposition into a finite set of sinusoids at the frequencies given by Eq. (39). Equivalently, because discrete complex exponentials form an orthonormal basis for all periodic sequences, the set of frequencies that uniquely represent the signal is entirely determined by the sequence length $N$ and the sampling rate $F_{s}$.
The periodicity of the time-domain sequence follows from the mathematical duality between sampling and periodic repetition in the time and frequency domains. Specifically, sampling the DTFT at $N$ equally spaced points implicitly converts the finite-length sequence into a periodic one, satisfying $x_{\text{in}}[N\cdot n]=x_{\text{in}}[N\cdot(n+1)]$. Finally, because both factors in the summation of Eq. (38) are periodic modulo $N$, the resulting sampled spectrum can likewise be repeated across the discrete frequency axis.
@@ V Discrete-Time LTI Systems and the Z-Domain
The role played by complex exponentials as eigenfunctions and modal decompositions in CT LTI systems is assumed by geometric sequences in the discrete-time domain, as this section will detail. While the transfer function of a continuous‑time system, Eq. 2, is derived via the Laplace transform, discrete‑time systems are naturally characterized using the Z‑transform, leading to a transfer function of the form polynomial ratio of the type:
@@
In the DT domain, the infinitesimal operations of differentiation and integration are replaced by finite differences and accumulation respectively, providing a simple algebraic relation between the samples in the sequence. In the Z‑domain as well, transfer functions offer clear insight into frequency‑domain behavior and impulse‑response characteristics, while simplifying the analysis of interactions between subsystems. These features are essential for enabling systematic, modular design of discrete‑time systems. As for the CT case, strict causality imposes a backward definition of the derivatives. Therefore the time-discrete derivatives of a sequence $x[k]=x(t)|_{t=k\cdot T_{s}}$ are defined as: